# A Statistical Framework of Watermarks for Large Language Models: Pivot, Detection Efficiency and Optimal Rules

### v3

@@ C.2 Key Lemma for Asymptotic Analysis
Let $\mu_{n}$ and $\mu$ be probability measures on ${\mathbb{R}}$ with CDFs given by $F_{n}$ and $F$ respectively. If $\mu_{n}$ weakly converges to $\mu$ and $F$is continuous, then $F_{n}$ converges uniformly to $F$ on ${\mathbb{R}}$.
@@ D.1 Simulation Setup
Hash collisions occur when the computed pseudorandom numbers are not truly i.i.d. It makes the way using $\hat{\gamma}_{n,\alpha}$ to control Type I error become invalid. This issue is well illustrated in Figure 1 of [22], where it is shown that the empirical false positive rates for common watermarks often exceed theoretical predictions. The primary reason for this discrepancy is that their pseudorandom numbers rely only on local information, i.e., $\zeta_{t}={\mathcal{A}}(w_{(t-m):(t-1)},{\mathtt{Key}})$ in our notation. Consequently, if there exists any $t\neq t^{\prime}$ such that $w_{(t-m):(t-1)}=w_{(t^{\prime}-m):(t^{\prime}-1)}$, then $\zeta_{t}=\zeta_{t^{\prime}}$, leading to the same pseudorandom numbers appearing multiple times. This violates the assumption of true i.i.d. nature and thus our working hypothesis. To mitigate this, [22] proposed heuristic methods to address such repetitions (see their Part C in Section III). By implementing these methods, they were able to improve Type I error control, as demonstrated in the rightmost panel of Figure 1.

### v4

@@ C.2 Key Lemma for Asymptotic Analysis
Let $\mu_{n}$ and $\mu$ be probability measures on $\mathbb{R}$} with CDFs given by $F_{n}$ and $F$ respectively. If $\mu_{n}$ weakly converges to $\mu$ and FFis continuous, then $F_{n}$ converges uniformly to $F$ on $\mathbb{R}$}.
@@ D.1 Simulation Setup
Hash collisions occur when the computed pseudorandom numbers are not truly i.i.d. It makes the way using $\hat{\gamma}_{n,\alpha}$ to control Type I error become invalid. This issue is well illustrated in Figure 1 of [22], where it is shown that the empirical false positive rates for common watermarks often exceed theoretical predictions. The primary reason for this discrepancy is that their pseudorandom numbers rely only on local information, that is, $\zeta_{t}={\mathcal{A}}(w_{(t-m):(t-1)},{\mathtt{Key}})$ in our notation. Consequently, if there exists any $t\neq t^{\prime}$ such that $w_{(t-m):(t-1)}=w_{(t^{\prime}-m):(t^{\prime}-1)}$, then $\zeta_{t}=\zeta_{t^{\prime}}$, leading to the same pseudorandom numbers appearing multiple times. This violates the assumption of true i.i.d. nature and thus our working hypothesis. To mitigate this, [22] proposed heuristic methods to address such repetitions (see their Part C in Section III). By implementing these methods, they were able to improve Type I error control, as demonstrated in the rightmost panel of Figure 1.